The electric field in a copper wire of cross sectional area 2 mm^2 carrying a current of 1 A

Grade:11

1 Answers

Arun
5 years ago
Dear Sreyas
 
A = 2 * 10^-6 m^2,
l = 1 A f = 1.7 * 10^-8 Ω-m
E = ?
R = fℓ/A = 1.7 * 10^-8 * ℓ/2 * 10^-6
V = IR = 1 * 1.7 * 10^-8 * ℓ/2 * 10^-6
E = dV/dL = V/l = 1.7 * 10^-8 * ℓ/2 * 10^-6 ℓ = 1.7/2 * 10^-2V/m = 8.5 mV/m.
 
Regards
Arun (askIITians forum expert)